A point moves along a circle having a radius 20 cm with a constant tangential acceleration


5 cm/s2. How much time is needed after motion begins for the normal acceleration of the point to be


equal to tangential acceleration ?


Grade:11

3 Answers

ankitesh gupta
11 years ago

2 SECONDS

mohit yadav
11 years ago

so,

 V2/R=5

 AND R=20cm

  so V=10cm/s

        and acceleration is dV/dt

      5=dv/dt

integrating we get time is 2seconds

Akash Kumar Dutta
11 years ago

v^2/r=5   (since normal acc=v^2/r=tangential acc.=dv/dt=5)

v^2=5.r=5.20=100.

v=10 cm/s

but dv/dt=5

dv=5dt

v=5t

10=5t

t=2 secs.(ANS)